OpenStudy (anonymous):

hey how do i put the complex number -6 into standard form

OpenStudy (anonymous):

Assuming that is the real part of the number then it would be -6+0i

OpenStudy (anonymous):

do you mean \[\sqrt{-6}\]

OpenStudy (anonymous):

naww just -6

OpenStudy (anonymous):

is so its \[\sqrt{-1}\sqrt{6}\]

OpenStudy (anonymous):

If all you're given is -6 and told to write it as a complex number then it's -6+0i

OpenStudy (anonymous):

u want it in standard form tho -6 is already a complex number

OpenStudy (anonymous):

If -6 is a complex number then -6+0i would be standard form.